Find the gradient of a line which is perpendicular to line y = -2x + 1
Explanation please. Thanks

Answer by Cromlix(4381)

Hi there,
y = -2x + 1
Gradient = -2
Lines that are perpendicular to
one another have gradients that
multiply together to give -1
m1 x m2 = -1
-2 x m2 = -1
m2 = -1/-2
m2 = 1/2
The gradient of the line
perpendicular to y = -2x + 1
is 1/2
hope this helps :-)